# Permutation groups, partition lattices and block structures

### v2

@@ Abstract
Let $G$ be a transitive permutation group on $\Omega$. The $G$-invariant partitions form a sublattice of the lattice of all partitions of $\Omega$, having the further property that all its elements are uniform (that is, have all parts of the same size). If, in addition, all the equivalence relations defining the partitions commute, then the relations form an orthogonal block structure, a concept from statistics; in this case the lattice is modular. If it is distributive, then we have a poset block structure, whose automorphism group is a generalised wreath product. We examine permutation groups with these properties, which we call the OB property and PB property respectively, and in particular investigate when direct and wreath products of groups with these properties also have these properties.
A famous theorem on permutation groups asserts that a transitive imprimitive group $G$ is embeddable in the wreath product of two factors obtained from the group (the group induced on a block by its setwise stabiliser, and the group induced on the set of blocks by $G$). We extend this theorem to groups with the PB property, embeddng them into generalised wreath products. We show that the map from posets to generalised wreath products preserves intersections and inclusions.
@@ 1 Introduction
Let $G$ be a transitive permutation group on a finite set $\Omega$. Then the $G$-invariant partitions of $\Omega$ form a sublattice of the lattice of all partitions of $\Omega$ (ordered by refinement). The $G$-invariant partitions have the additional property that they are uniform (all parts have the same size).
In this paper we are primarily interested in the class of permutation groups for which the equivalence relations corresponding to the $G$-invariant partitions commute pairwise. (We will see in Section 5.1 that, at least among transitive groups of small degree, the vast majority do satisfy this condition; for example, $1886$ of the $1954$ transitive groups of degree $16$ do so.) Then the lattice of partitions which they form is called an orthogonal block structure, for short an OBS. This property can also be defined by saying that the subgroups containing a point stabiliser $G_{\alpha}$ commute pairwise. This implies that the lattice satisfies the modular law. It turns out that this property of a partition lattice was introduced, in the context of statistical design, by several different statisticians: see Section 3.
@@ 2.2 Lattices
We are required to prove that $\Phi\preccurlyeq\Psi$ implies Φ(ΞΨ)=($\Phi\vee(\Xi\wedge\Psi)=(\Phi\vee\Xi)\wedge\Psi$. In Figure 1, the dots represent points in $\Omega$. Each edge is labelled by a partition of $\Omega$. If an edge labelled $\Phi$ joins points $\alpha$ and $\beta$, this means that $\alpha$ and $\beta$ are in the same part of $\Phi$; and similarly for $\Xi$ and $\Psi$.
@@
A finite distributive lattice $L$ is isomorphic to the lattice of down-sets in a partially ordered set $M$. We can take $M$ to be the set of join-indecomposable elements of $L$ (elements $m$ satisfying $m=m_{1}\vee m_{2}$ implies $m=m_{1}$ or $m=m_{2}$).
A proof of this theorem is in [11, p. 192]. We sometimes abbreviate “join-indecomposable” to JI.
@@
The proof of this theorem can be found in [15, p. 134].
@@ 2.3 Orthogonal block structures
The set =$\mathcal{B}=\{E,U\}$ is an orthogonal block structure, which we call trivial.
@@
Take a complete set of $q-1$ mutually orthogonal Latin squares of order $q$. Take $\Omega$ to be the set of cells of the square; as well as the partitions $E$ and $U$, take the partitions into rows, columns, and letters of each of the squares. We obtain an orthogonal block structure. Since every pair of cells are either in the same row or column or carry the same letter in one of the squares, applying the above construction to the relation $U$ gives the empty relation. So the association scheme has $q+1$ classes apart from the diagonal.
@@
We conclude with two remarks on association schemes.
@@ 2.4 Crossing and nesting
If $R_{1}$ and $R_{3}$ are in $\mathcal{B}_{1}$ and $R_{2}$ and $R_{4}$ are in $\mathcal{B}_{2}$ then $(R_{1}\circ R_{3})\times(R_{2}\circ R_{4})=(R_{1}\times R_{2})\circ(R_{3}\times R_{4})$. Therefore, if every two equivalence relations in $\mathcal{B}_{1}$ commute and every two equivalence relations in $\mathcal{B}_{2}$ commute, then the same is true for every two equivalence relations in each of $\mathcal{P}_{1}\times\mathcal{P}_{2}$, $\mathcal{P}_{1}/\mathcal{P}_{2}$ and $\mathcal{P}_{2}/\mathcal{P}_{1}$.
For permutation group theorists, note the similarities between crossing and nesting on one hand, and direct product (with product action) and wreath product (with imprimitive action) on the other. Statisticians call the results of crossing and nesting trivial OBSs row-column structures and block structures respectively.
@@ 2.5 Poset block structures
where $\alpha_{i},\beta_{i}\in\Omega_{i}$ for all $m_{i}\in M$. Then $R_{D}$ is an equivalence relation on $\Omega$, and we let $\Pi_{D}$ be the corresponding partition.
@@
For every $i\in\{1,\ldots,N\}$, let $A(i)$ denote the set $\{j\in\{1,\ldots,N\}\,:\,m_{i}\sqsubset m_{j}\}$ and $A[i]$ the set $\{j\in\{1,\ldots,N\}\,:\,m_{i}\sqsubseteq m_{j}\}$. Similarly, let $D(i)$ denote the set $\{j\in\{1,\ldots,N\}\,:\,m_{j}\sqsubset m_{i}\}$ and $D[i]$ the set $\{j\in\{1,\ldots,N\}\,:\,m_{j}\sqsubseteq m_{i}\}$. (Mnemonic: $A={}$‘ancestor’, $D={}$‘descendant’.)
@@ 2.6 Generalised wreath products
The generalised wreath product $G$ of the groups $G(m_{1}),\ldots,G(m_{N})$ over the poset $M$ is the group $\prod_{i=1}^{N}F_{i}$, and it acts on $\Omega$ in the following way: if $\omega=(\omega_{1},\ldots,\omega_{N})\in\Omega$ and $f=\prod_{i=1}^{N}f_{i}\in G$, then
@@ 3.1 Fisher and Yates at Rothamsted
[…]
Split-plot design $\underline{\underline{b}}$$\underline{\underline{k}}$$\underline{\underline{s}}$ $U$$b$$bk$$bks$
@@
Two Latin squares with subplots $\underline{\underline{4}}$$\underline{\underline{4}}$$\underline{\underline{2}}$$\underline{\underline{2}}$ $U$$1$squares2222plots32323232subplots646464648888rows8888columns
[…]
Four Latin squares of order five $\underline{\underline{4}}$$\underline{\underline{5}}$$\underline{\underline{5}}$ $U$$1$squares444420202020rows20202020columns100100100100plots
@@ 3.3 Statisticians at Iowa State University
With hindsight, it seems that they were trying to define poset block structures, but they managed to confuse the poset $M$ of coordinates with the lattice of partitions. They denoted the universal partition $U$ by $\mu$, and the equality partition $E$ by $\varepsilon$. They used complicated formulae, called symbolic representations, to explain the partial order $M$, but then included $\mu$ and $\varepsilon$ in the corresponding Hasse diagram, which they called the structure diagram. They dealt with all posets of size at most four, and showed $16$ of the $63$ posets of size five.
@@
$S\colon RC$ $\mu$$S$$C$$R$$\varepsilon$ $U$$S$$R$$C$$E$
$S\colon(R)(C\colon L)$ $\mu$$S$$R$$L$$C$$\varepsilon$ $S$$U$$C$$L$$E$$C\wedge R$$R$
$(S\colon Q)$ $(P)$ and $(SP\colon R)$ $\mu$$\varepsilon$$P$$R$$S$$Q$ $U$$P$$S$$Q$$P\wedge S$$R$$P\wedge Q$$E$
@@ 3.4 Unifying the theory
They also explained that most of the theory extends to what we now call an orthogonal block structure, where the lattice is modular but not necessarily distributive. Figure 6 shows the corresponding Hasse diagrams in their two examples. In the one on the left, the non-trivial partitions form the rows, columns, Latin letters and Greek letters of a pair of mutually orthogonal Latin squares, so the underlying set has size $n^{2}$ with $n\notin\{1,2,6\}$. One way of achieving the one on the right is to use some carefully chosen subgroups of the elementary abelian group of order $16$.
@@ 4.1 Introduction to OB groups
We observe that, for a given point $\alpha\in\Omega$, there is a natural order-preserving bijection between $G$-invariant partitions of $\Omega$ and subgroups of $G$ containing $G_{\alpha}$: if $G_{\alpha}\leq H\leq G$, then $\alpha H$ is a part of a $G$-invariant partition; in the other direction, if $\Pi$ is a $G$-invariant partition, the corresponding subgroup is the setwise stabiliser of the part of $\Pi$ containing $\alpha$. If $\Pi_{1}$ and $\Pi_{2}$ correspond to $H$ and $K$, then $\Pi_{1}\wedge\Pi_{2}$ corresponds to $H\cap K$, and $\Pi_{1}\vee\Pi_{2}$ corresponds to $\langle H,K\rangle$. (The result for join is in [1], and for meet [16, Theorem 1.5A].
@@
Conversely, suppose that $\Pi_{1}$ and $\Pi_{2}$ are the $G$-invariant partitions corresponding to $H$ and $K$, and that $\Pi_{1}\circ\Pi_{2}=\Pi_{1}\vee\Pi_{2}$. In particular, this holds for the part containing $\alpha$. So any point in this part can be reached from $\alpha$ by first moving to a point $\beta$ in the same part of $\Pi_{1}$, then to a point $\gamma$ in the same part of $\Pi_{2}$ as $\beta$. Since the stabiliser of the part of $\Pi_{1}$ containing $\alpha$ is $H$, we have $\beta=\alpha h$ for some $h\in H$. Then the part of $\Pi_{2}$ containing $\beta$ is obtained by mapping the part containing $\alpha$ by $h$, so its stabiliser is $K^{h}$; so $\gamma=\beta h^{-1}kh$ for some $k\in K$. Thus $\gamma=\alpha kh$. We conclude that the part of $\Pi_{1}\vee\Pi_{2}$ containing $\alpha$ is $\alpha KH$. Because the partitions commute, this part is also equal to $\alpha HK$. We conclude that $HK=KH$.  $\Box$
@@
    Suppose that $n=p^{2}$, with $p$ prime. Any non-trivial $G$-invariant partition has $p$ parts of size $p$; the meet and join of two such partitions are trivial. So $G$ has the OB property. If there are at most two such partitions, then the OBS preserved by $G$ is obtained by crossing or nesting two trivial structures of size $p$, and so $G$ is embedded in the direct or wreath product of two transitive groups of degree $p$.
    Suppose that $G$ has more than two non-trivial invariant partitions. Then the Sylow $p$-subgroup $P$ of $G$ is regular and elementary abelian, so $G$ preserves a Latin square which is the Cayley table of $C_{p}$. Then $P$ fixes $p+1$ non-trivial partitions, forming the parallel classes of lines in the affine plane over the field of $p$ elements; all these partitions are fixed by $G$. Thus $G$ is embedded in the affine group $\mathop{\mathrm{AGL}}(2,p)$.

### v3

@@ Abstract
Let $G$ be a finite transitive permutation group on $\Omega$. The $G$-invariant partitions form a sublattice of the lattice of all partitions of $\Omega$, having the further property that all its elements are uniform (that is, have all parts of the same size). If, in addition, all the equivalence relations defining the partitions commute, then the relations form an orthogonal block structure, a concept from statistics; in this case the lattice is modular. If it is distributive, then we have a poset block structure, whose automorphism group is a generalised wreath product. We examine permutation groups with these properties, which we call the OB property and PB property respectively, and in particular investigate when direct and wreath products of groups with these properties also have these properties.
A famous theorem on permutation groups asserts that a transitive imprimitive group $G$ is embeddable in the wreath product of two factors obtained from the group (the group induced on a block by its setwise stabiliser, and the group induced on the set of blocks by $G$). We extend this theorem to groups with the PB property, embedding them into generalised wreath products. We show that the map from posets to generalised wreath products preserves intersections and inclusions.
@@ 1 Introduction
Let $G$ be a finite transitive permutation group on a finite set $\Omega$. Then the $G$-invariant partitions of $\Omega$ form a sublattice of the lattice of all partitions of $\Omega$ (ordered by refinement). The $G$-invariant partitions have the additional property that they are uniform (all parts have the same size).
In this paper all permutation groups will be finite and we are primarily interested in the class of permutation groups for which the equivalence relations corresponding to the $G$-invariant partitions commute pairwise. (We will see in Section 5.1 that, at least among transitive groups of small degree, the vast majority do satisfy this condition; for example, $1886$ of the $1954$ transitive groups of degree $16$ do so.) Then the lattice of partitions which they form is called an orthogonal block structure, for short an OBS. This property can also be defined by saying that the subgroups containing a point stabiliser $G_{\alpha}$ commute pairwise. This implies that the lattice satisfies the modular law. It turns out that this property of a partition lattice was introduced, in the context of statistical design, by several different statisticians: see Section 3.
@@
Acknowledgements. We are grateful to Michael Kinyon for drawing our attention to the paper [38]. We thank the anonymous referees for a careful reading of the paper and all the useful comments and suggestions that have greatly improved the paper. The first author thanks the School of Mathematics and Statistics of the University of St Andrews for supporting part of this work with an undergraduate StARIS scholarship. The rest of the work was undertaken during the first author’s doctoral studies at the University of Bristol and she thanks the Heilbronn Institute for Mathematical Research for the financial support.
@@ 2.2 Lattices
We are required to prove that $\Phi\preccurlyeq\Psi$ implies $\Phi\vee(\Xi\wedge\Psi)=(\Phi\vee\Xi)\wedge\Psi$. In Figure 1, the dots represent points in $\Omega$. Each edge is labelled by a partition of $\Omega$. If an edge labelled $\Phi$ joins points $\alpha$ and $\beta$, this means that $\alpha$ and $\beta$ are in the same part of $\Phi$; and similarly for $\Xi$ and $\Psi$.
@@
Proposition 2.3 is a standard result in lattice theory and appears in [16] as Lemmas 4.2 and 4.3.
@@
A finite distributive lattice $L$ is isomorphic to the lattice of down-sets in a partially ordered set $M$. We can take $M$ to be the set of join-indecomposable elements of $L$ (elements $m$ for which $m=m_{1}\vee m_{2}$ implies $m=m_{1}$ or $m=m_{2}$).
A proof of this theorem is in [12, p. 192]. We sometimes abbreviate “join-indecomposable” to JI.
@@
The proof of this theorem can be found in [16, p. 134].
@@ 2.3 Orthogonal block structures
The set $\mathcal{B}=\{E,U\}$ is an orthogonal block structure, which we call trivial.
@@
Recall that two Latin squares $A=(a_{ij})$ and $B=(b_{ij})$ over alphabets $\mathcal{A}$ and $\mathcal{B}$ are orthogonal if for every pair $(a,\beta)\in\mathcal{A}\times\mathcal{B}$ there exists a unique pair $(i,j)$ such that $a_{ij}=a$ and $b_{ij}=\beta$. Note that it is a convention going back to Euler that the two alphabets are the Latin and Greek letters respectively when dealing with two orthogonal Latin squares.
Take a complete set of q1q-1 mutually orthogonal Latin squares of order $q$. Take $\Omega$ to be the set of cells of the square and let $\mathcal{B}$ be the set containing the partitions $E$ and $U$, and the partitions into rows, columns, and letters of each of the squares. We claim that $\mathcal{B}$ is an orthogonal block structure.
Conditions (a) and (c) are straightforward to verify and (b) holds by construction, so we will only show that the partitions commute. Let $\Pi_{1},\Pi_{2}\in\mathcal{B}$. If $\Pi_{1}$ and $\Pi_{2}$ are the row and the column partitions, then they commute because any cell can be reached by first moving along a row and then a column or vice versa.
If $\Pi_{1}$ is the row partition and $\Pi_{2}$ is a letter partition, then we note that each letter appears in every row, and so if $x_{1}$ and $x_{2}$ are two cells in rows $r_{1}$ and $r_{2}$ respectively containing the letters $a$ and $b$ in the Latin square corresponding to $\Pi_{2}$, then we can reach $x_{2}$ from $x_{1}$ either by first moving to the cell in $r_{1}$ containing the letter $b$ and then to $x_{2}$, or by first moving to the cell containing the letter $a$ in $r_{2}$ and then moving along $r_{2}$ to $x_{2}$. Therefore $\Pi_{1}$ and $\Pi_{2}$ commute. An entirely similar argument shows that $\Pi_{1}$ and $\Pi_{2}$ commute if $\Pi_{1}$ is the column partition.
Finally, suppose that $\Pi_{1}$ and $\Pi_{2}$ are both letter partitions and let $L_{1}$ and $L_{2}$ be the Latin squares corresponding to $\Pi_{1}$ and $\Pi_{2}$ respectively. Moreover, let $x_{1}$ and $x_{2}$ be cells that contain the letters $a$ and $b$ in $L_{1}$ and the letters $\alpha$ and $\beta$ in $L_{2}$. Orthogonality ensures that there is exactly one cell $x_{3}$ which contains the letter $a$ in the $L_{1}$ and the letter $\beta$ in $L_{2}$, so we can move from $x_{1}$ to $x_{2}$ via $x_{3}$. Similarly, there exists a unique cell $x_{4}$ that contains the letter $\alpha$ in $L_{2}$ and the letter $b$ in $L_{1}$, so we can move from $x_{1}$ to $x_{2}$ via $x_{4}$, and so $\Pi_{1}\circ\Pi_{2}=\Pi_{2}\circ\Pi_{1}$.
Since every pair of cells are either in the same row or column or carry the same letter in one of the squares, constructing the association scheme obtains the empty relation from the universal partition $U$. So the association scheme has $q+1$ classes apart from the diagonal.
@@
Let us now record some remarks on association schemes.
@@
It was pointed out to us by the anonymous referee that partition lattices whose equivalence relations commute were independently introduced by G.-C. Rota and his students and collaborators, who called them linear lattices. Details can be found in [19], where it is explained how they arise in lattice theory and are connected with logic. The authors agree with the referee’s view that Rota would have enjoyed the connection with OBSs.
We conclude this section by stating a result by Dubreil and Dubreil-Jacotin [18] that gives an alternative characterisation of partition lattices whose equivalence relations commute. We will say that two partitions $\Pi_{1}$ and $\Pi_{2}$ are independent, if for every $P_{1}\in\Pi_{1}$ and $P_{2}\in\Pi_{2}$ we have $P_{1}\cap P_{2}\neq\emptyset$.
Proposition 2.6
Two partitions $\Pi_{1}$ and $\Pi_{2}$ commute if and only if for every part $P\in\Pi_{1}\vee\Pi_{2}$, the restrictions $\Pi_{1}|_{P}$ and $\Pi_{2}|_{P}$ are independent.
@@ 2.4 Crossing and nesting
If $R_{1}$ and $R_{3}$ are in $\mathcal{B}_{1}$ and $R_{2}$ and $R_{4}$ are in $\mathcal{B}_{2}$ then $(R_{1}\circ R_{3})\times(R_{2}\circ R_{4})=(R_{1}\times R_{2})\circ(R_{3}\times R_{4})$. Therefore, since every two equivalence relations in $\mathcal{B}_{1}$ commute and every two equivalence relations in $\mathcal{B}_{2}$ commute, then the same is true for every two equivalence relations in each of $\mathcal{P}_{1}\times\mathcal{P}_{2}$, $\mathcal{P}_{1}/\mathcal{P}_{2}$ and $\mathcal{P}_{2}/\mathcal{P}_{1}$.
Let $G\leq{\rm Sym}(\Gamma)$ and $H\leq{\rm Sym}(\Delta)$ be transitive permutation groups. Recall that the product action of the direct product $G\times H$ on $\Gamma\times\Delta$ is the action defined by $(\gamma,\delta)(g,h)=(\gamma g,\delta h)$ for all $(\gamma,\delta)\in\Gamma\times\Delta$ and $(g,h)\in G\times H$. For permutation group theorists, note the similarities between crossing and nesting on one hand, and direct product (with product action) and wreath product (with imprimitive action) on the other. Statisticians call the results of crossing and nesting trivial OBSs row-column structures and block structures respectively.
@@ 2.5 Poset block structures
where $\alpha_{i},\beta_{i}\in\Omega_{i}$ for all $m_{i}\in M$. Then $R_{D}$ is an equivalence relation on $\Omega$, and we let $\Pi_{D}$ be the corresponding partition. (The appearance of the poset $M$ explains the name poset block structures.)
@@
For every $i\in\{1,\ldots,N\}$, let $A(i)$ denote the set $\{j\in\{1,\ldots,N\}\,:\,m_{i}\sqsubset m_{j}\}$ and $A[i]$ the set $\{j\in\{1,\ldots,N\}\,:\,m_{i}\sqsubseteq m_{j}\}$. Similarly, let $D(i)$ denote the set $\{j\in\{1,\ldots,N\}\,:\,m_{j}\sqsubset m_{i}\}$ and $D[i]$ the set $\{j\in\{1,\ldots,N\}\,:\,m_{j}\sqsubseteq m_{i}\}$. (Mnemonic: A=A={}‘ancestor’, D=D={}‘descendant’.)
@@ 2.6 Generalised wreath products
The generalised wreath product $G$ of the groups $G(m_{1}),\ldots,G(m_{N})$ over the poset $M$ is the group $\prod_{i=1}^{N}F_{i}$, and it acts on $\Omega$ in the following way: if $\omega=(\omega_{1},\ldots,\omega_{N})\in\Omega$ and $f=(f_{1},\ldots,f_{N})\in G$, then
@@ 3.1 Fisher and Yates at Rothamsted
Block design $\underline{\underline{k}}$$\underline{\underline{b}}$ UU11blocksbbplotsbkbk
@@
Split-plot design $\underline{\underline{b}}$$\underline{\underline{k}}$$\underline{\underline{s}}$ UUblocksplotssubplots11bbbkbkbksbks
@@
Two Latin squares with subplots $\underline{\underline{4}}$$\underline{\underline{4}}$$\underline{\underline{2}}$$\underline{\underline{2}}$ UU11squares22plots3232subplots646488rows88columns
Splitting rows of a rectangle $\underline{\underline{r}}$$\underline{\underline{2}}$$\underline{\underline{c}}$ UU11rrrows2r2rsubrows2rc2rcsubplotsrcrcplotscolumnscc
Four Latin squares of order five $\underline{\underline{4}}$$\underline{\underline{5}}$$\underline{\underline{5}}$ UU11squares442020rows2020columns100100plots
@@ 3.3 Statisticians at Iowa State University
With hindsight, it seems that they were trying to define poset block structures, but they managed to confuse the poset $M$ of coordinates with the lattice of partitions. They denoted the universal partition $U$ by $\mu$, and the equality partition $E$ by $\varepsilon$. They used complicated formulae, called symbolic representations, to explain the poset $M$, but then included $\mu$ and $\varepsilon$ in the corresponding Hasse diagram, which they called the structure diagram. They dealt with all posets of size at most four, and showed $16$ of the $63$ posets of size five.
@@
[…]
@@ 3.4 Unifying the theory
They also explained that most of the theory extends to what we now call an orthogonal block structure, where the lattice is modular but not necessarily distributive. Figure 6 shows the corresponding Hasse diagrams in their two examples. In the one on the left, the non-trivial partitions form the rows, columns, and the Greek and Latin letter partitions of a pair of mutually orthogonal Latin squares. Note that the underlying set has size $n^{2}$ with $n\notin\{1,2,6\}$, since it is well known that there exists a pair of two mutually orthogonal Latin squares of order $n$ if and only if $n\notin\{1,2,6\}$. One way of achieving the one on the right is to use some carefully chosen subgroups of the elementary abelian group of order $16$.
@@ 4.1 Introduction to OB groups
We observe that, for a given point $\alpha\in\Omega$, there is a natural order-preserving bijection between $G$-invariant partitions of $\Omega$ and subgroups of $G$ containing $G_{\alpha}$: if $G_{\alpha}\leq H\leq G$, then $\alpha H$ is a part of a $G$-invariant partition; in the other direction, if $\Pi$ is a $G$-invariant partition, the corresponding subgroup is the setwise stabiliser of the part of $\Pi$ containing $\alpha$. If $\Pi_{1}$ and $\Pi_{2}$ correspond to $H$ and $K$, then $\Pi_{1}\wedge\Pi_{2}$ corresponds to $H\cap K$, and $\Pi_{1}\vee\Pi_{2}$ corresponds to $\langle H,K\rangle$. (The result for join is in [1], and for meet [17, Theorem 1.5A].
@@
Conversely, suppose that $\Pi_{1}$ and $\Pi_{2}$ are the $G$-invariant partitions corresponding to $H$ and $K$, and that $\Pi_{1}\circ\Pi_{2}=\Pi_{1}\vee\Pi_{2}$. In particular, this holds for the part containing $\alpha$. So any point in this part can be reached from $\alpha$ by first moving to a point $\beta$ in the same part of $\Pi_{1}$, then to a point $\gamma$ in the same part of $\Pi_{2}$ as $\beta$. Since the stabiliser of the part of $\Pi_{1}$ containing $\alpha$ is $H$, we have $\beta=\alpha h$ for some $h\in H$. Then the part of $\Pi_{2}$ containing $\beta$ is obtained by mapping the part containing $\alpha$ by $h$, so its stabiliser is $K^{h}$; so $\gamma=\beta h^{-1}kh$ for some $k\in K$. Thus $\gamma=\alpha kh$. We conclude that the part of $\Pi_{1}\vee\Pi_{2}$ containing $\alpha$ is $\alpha KH$. Because the partitions commute, this part is also equal to $\alpha HK$. We now claim that this implies that $HK=KH$.
Let $g_{1}\in HK$. Then $\alpha g_{1}\in\alpha HK=\alpha KH$, so there exists some $g_{2}\in KH$ such that $\alpha g_{1}=\alpha g_{2}$. It follows that $\alpha g_{1}g_{2}^{-1}=\alpha$, and so $g_{1}g_{2}^{-1}\in G_{\alpha}$, which in turn implies that $g_{1}\in G_{\alpha}g_{2}\subseteq G_{\alpha}KH$. But since $G_{\alpha}\leq H$ and $G_{\alpha}\leq K$, we have $G_{\alpha}KH=KH$, and so $g_{1}\in KH$. Therefore $HK\subseteq KH$. By symmetry, we also get $KH\subseteq HK$, and thus $HK=KH$, as claimed.   $\Box$
@@
    Suppose that $n=p^{2}$ for some prime $p$ and that there are more than two non-trivial $G$-invariant partitions. Then each such partition has $p$ parts of size $p$, and any two have meet $E$ and join $U$. Thus any three of these partitions give the structure of a Latin square $L$ to the underlying set $\Omega$.
    Let $P$ be a Sylow $p$-subgroup of $G$. The stabiliser in $P$ of a part of a non-trivial $G$-invariant partition $\Pi_{1}$ has index $p$ in $G$ by the Orbit-Stabiliser Theorem, and fixes all parts of $\Pi_{1}$. If $\Pi_{2}$ is another such partition, then the stabiliser in $P$ of a part in each of $\Pi_{1}$ and $\Pi_{2}$ fixes all parts of $\Pi_{1}\wedge\Pi_{2}$, that is, it is the identity. So $|P|=p^{2}$; and $P$, having more than one subgroup of index $p$, is the elementary abelian group. Since $P$ induces a cyclic group $C_{p}$ of order $p$ on each of the sets of rows, columns and letters of $L$, we see that $L$ is the Cayley table of $C_{p}$.
    The automorphism group of this Latin square is $(C_{p}\times C_{p})\colon C_{p-1}$. So $G$ is contained in this group. But $C_{p}\times C_{p}$ has $p+1$ non-trivial invariant partitions (corresponding to its $p+1$ subgroups of order $p$) and $C_{p-1}$ fixes just two of these partitions and permutes the other p1p-1 regularly. Since $G$ fixes at least three partitions, we conclude that $G=C_{p}\times C_{p}$. The $p+1$ non-trivial $G$-invariant partitions together with $U$ and $E$ form an orthogonal block structure which is not a poset block structure.